OpenStudy (anonymous):

as x > infinite what is the limit of ( 1/(a/x) )^ bx

OpenStudy (anonymous):

\[\lim_{x\to\infty}\left(\frac{1}{a/x}\right)^{bx}~~?\]

OpenStudy (anonymous):

yesss

OpenStudy (anonymous):

Depends on \(b\).

OpenStudy (anonymous):

it wants the limit usin l'hopitals tho

OpenStudy (anonymous):

a should be positive. otherwise you can't compute all powers (i think).

OpenStudy (anonymous):

if b > 0 then you basically have infty^infty = infty

OpenStudy (anonymous):

if b = 0 you have 1

OpenStudy (anonymous):

if b < 0 let c = -b > 0. You have \(\lim_{x\to\infty}\left(\frac{a}{x}\right)^{cx}\). The base alone approaches 0, and an increasing exponent won't stop it.